# Tabby: Tabular Data Synthesis with Language Models

### v1

@@ 3.2 Investigating the Choice of Base Model
Claim 2: The Tabby architecture modification allows smaller LLMs to achieve synthetic data fidelity similar or better than LLMs with higher parameter counts.
Comparisons: We compare tabular synthesis quality across LLMs of varying sizes. We consider seven LLMs, listed in Table 3 evaluating Non-Tabby and MH versions of each. Each model is plain-trained under the conditions provided in Section 3.0, then sampled 500 times. Results are averaged across two runs. Llama models use LoRA [10] on all linear transformer layers, with the LM head fully fine-tuned.
@@
We find that Tabby improves MLE or maintains upper-bound MLE for models. Although higher-parameter models are generally correlated with greater generative abilities, Figure 5 demonstrates that this is not always the case: Interestingly, we find that the Llama models (1.2B and 8B parameters each), have lower average MLE than all smaller models. Meanwhile, Tabby offers favorable performance improvements relative to the scaling curve and allows even small models to better outperform large, resource-intensive models.
@@ 3.3 Extending Tabby Beyond Tabular Data to General Structured Modalities
Claim 3: Tabby architecture modifications may also be applied to other structured data, resulting in higher-quality synthetic data in these modalities as well.
Comparisons: We plain-train non-Tabby and Tabby MH models on a JSON dataset of patients being evaluated for Glaucoma [15]. Each datapoint has 10 features, organized in 3 groups: a group of 7 columns representing qualitative aspects of the optic nerve, a group of 2 columns corresponding to measurements between the optic nerve and eye, and a standalone feature for the diagnosis (examples in Box C). The binary classification target is inside the first group and assesses whether the optic nerve is thinning.
As with tabular datasets in Section 3.1 and 3.2, we train downstream classifiers to predict the target variable and then present the resulting MLE.
Additionally, we consider the discrimination metric: given a dataset containing equal numbers of real and synthetic samples, we measure the accuracy of a discrimination classifier that is trained to distinguish real from synthetic data. Because accuracy would indicate that the classifier is fully unable to distinguish real from synthetic, we report the accuracy’s distance from in Table 4 so that lower scores indicate higher-quality synthesis.
@@
| Original (Upper Bound) | ||
@@
Results: Table 4 demonstrates that Tabby MH improves MLE by , to parity with real data. Additionally, Tabby MH’s lower discrimination score signifies that this model’s samples are more realistic than the non-Tabby samples.
@@ 3.4 Tracking the Adaptation to Individual Columns
We address our final claim by examining Tabby’s progress while fine-tuning on tabular data.
@@ 3.5 Discussion
We find that Tabby models synthesize high-quality data in a variety of settings. In particular, Plain-trained Tabby MH consistently outperforms all prior LLM-based approaches and is comparable to or better than Tab-DDPM in most settings, despite Tabby enjoying greater flexibility under fewer assumptions. The Tabby architecture modification allows LLMs to better model univariate column distributions, as well as multivariate relationships across columns. This is evident across model sizes, shown in Section 3.2.
Unusually, we find that the baseline Plain training technique with Distilled-GPT2 achieves near-optimal MLE on several datasets. The high performance of the Plain training technique compared to prior LLM works on tabular synthesis, which are more complex, is surprising. As of this writing, the Adult, House and Diabetes datasets have become quite prevalent for tabular synthesis evaluation. We hope that future research will begin to focus on more challenging datasets as well as extensions into other structured modalities.
@@ 4 Related Work
Tabular data has played a central role in machine learning since the field’s early days. In particular, decision trees and relatives [29] are well-adapted to table classification or regression. Table synthesis, meanwhile, is frequently overlooked in favor of image and text synthesis–though still possesses a growing body of work.
Classical synthesis: Classic machine learning models, such as Bayesian networks, may be used to synthesize tabular data [24, 35]. However, these approaches are limited in the data types and distributions that may be represented.
Generative Adversarial Networks (GANs): Many tabular synthesis methods rely on GANs [8, 33], but have encountered several limitations. In particular, distributions of ordinal columns are frequently imbalanced, leading GANs to undesirable phenomena such as mode collapse. Continuous columns may possess multiple modes and complex interactions with the other columns which GANs also struggle to capture.
Diffusion Models: Tab-DDPM [12], a current state-of-the-art table synthesis approach, is based on the diffusion model architecture. Tab-DDPM shows top performance on many standard tabular metrics and is a highly reliable approach for certain applications. Unfortunately, this performance is achieved with strong assumptions on the nature of the data space–for instance, numeric target variables may only assume integer values (see Figure 4) and diffusion models are unable to model non-categorical string columns such as addresses or telephone numbers. The ability to reach comparable performance to Tab-DDPM with fewer assumptions is as an area of active research.
LLMs: A small body of work has sought to apply LLMs’ demonstrated abilities of modeling complex relationships to tabular data. The landmark work in this area, GReaT [3], details methods to convert tabular data into a sentence format which may be input to LLMs, then proposes a training technique of “shuffling" the order in which columns occur for each row, which is reported to improve the modeling of inter-column dependencies.
Two notable works have built off of GReaT: TapTap [36] pretrains full or Distilled-GPT2 [22] on a variety of tabular data before fine-tuning on a downstream tabular synthesis task, while Tabula [38] explores methods of preprocessing the training data to be more-easily modeled by LLMs. Other LLM-based works have adapted these recent advances to relational tables [28], or used the emergent abilities of very large models such as GPT-4 to generate synthetic data using In-Context Learning in place of fine-tuning [26]. Many of these prior LLM-based works are training techniques and may be applied in concert with Tabby: we demonstrate this using GReaT, TapTap and Tabula in Section 3.
@@ 5 Conclusion
We introduce Tabby, an MOE-based architecture modification that allows LLMs generate realistic tabular data. Tabby reaches parity with real data in datasets, according to machine learning efficacy. Given the promising performance of Tabby, we hope to spur future work in this area of architecture modifications that allow LLMs to better fit to tabular data and structured data.
@@ Appendix A Terminology Glossary
GReaT [3]: The landmark work on fine-tuning pre-existing LLMs to synthesize tabular data by encoding datapoints as text. Similar to Plain training, but includes train-time complications such as shuffling the order in which columns are encoded and sample-time complications such as the inability to generate label values that do not occur in the training dataset. Discussed in-detail in Section 4.
GReaT+TapTap+Tabula (GTT): The combination of GReaT training plus Tabula encoding and TapTap [36] pre-training on tabular data (which is performed after the LLM is pre-trained on text data).
@@ Appendix B Additional dataset information
Descriptions of Datasets
Diabetes [11] contains medical information on female hospital patients, including age, number of pregnancies and skin thickness. Downstream models learn to predict whether a given patient suffers from diabetes. Apart from the label, all dataset columns are numerical, with some columns taking only integer values, while others are floats.
Travel [30] was collected by a travel agency wishing to predict customer churn. With the binary variable churn as the target, features include whether the traveler booked a hotel, frequent flyer status and traveler age. While most features are categorical, there are two numerical columns: traveler age and the number of times that the customer has used the travel agency in the past.
Adult [2] is a dataset commonly used to benchmark tabular classification algorithms. Each row contains basic information on one American adult, such as their age, years of education and marital status. For each adult, the downstream task is to predict whether their annual income is above or below . The features are a mix of categorical and numerical columns, with each numerical column taking only integer values.
Our first regression dataset is Abalone [17], which records the basic measurements of abalones, such weight and height. The target variable is the abalone’s age.
The Rainfall [34] dataset, while challenging to many LLM-based synthesis methods, contains only four columns which record historical weather data in Bangladesh. Its target variable is the amount of rainfall recorded, and the features are the year, month and weather station location.
House [19] is a standard regression dataset. Each row represents a block of houses in California during the 1990 census. The dataset records the number of households residing in the block, the block’s median building age, average number of bedrooms, and other basic information. The dataset’s target column is the block’s median house value, which is numerical and allows us to assess Tabby’s synthetic data in a regression task.
Glaucoma [15] dataset consists of JSON records describing ophthalmic patients under consideration for a glaucoma diagnosis. Each record contains various qualitative and quantitative information about the eye, as demonstrated by the examples in Box 1.
@@ Appendix D Details on Experimental Setup
Discrimination [21] quantifies the degree to which the generative model introduces spurious correlations or other patterns that differentiate synthetic from real data. Given the real training dataset and a synthetic dataset , we sample the same number of rows from each. Next, we train a random forest classifier to discriminate between real and synthetic examples. Highest-quality synthetic data will result in discrimination accuracy, indicating that is unable to distinguish between and . For this reason, our reported discrimination scores are calculated as the absolute difference between and the accuracy of discriminator . Accordingly, lower discrimination scores represent better performance.
Distance to Closest Record:
Distance to Closest Record (DCR) [13] quantifies the distance between each synthetic datapoint and its most-similar example in the training set . In addition to synthesis quality, this metric is an indication of the degree to which the model memorizes samples during training. Specifically, for each synthetic example , we compute its distance to every training point (using L0 distance for categorical columns and L1 distance for numerical columns) and take the smallest of these distances. The overall DCR is then reported as the average of these minimum distances across all synthetic examples in SS. Lower DCR is associated with high-quality synthesis, but a DCR score of 0 implies that most synthetic examples are merely copies of training dataset points memorized during training. As such, we consider the best DCR to be the lowest nonzero score.
@@ E.1 Additional Metrics for Main Results Tables
For the experiment presented in Section 3.1, we include the two additional metrics of discrimination and distance to closest record (DCR) in Tables 6 and 7 respectively. These metrics largely corroborate our findings in Section 3.1. In particular, Plain Tabby MH’s low DCR and discrimination scores indicate that this model’s synthetic data closely resembles that of real data. Additionally, the DCR scores are small but nonzero, which indicates that the model is generating novel datapoints rather than simply repeating datapoints memorized during training.
@@
| CTGAN | ||||||
| TVAE | ||||||
| Tab-DDPM | ||||||
| Plain Base DGPT2 | ||||||
| Plain Tabby MH DGPT2 | ||||||
| GReaT Base DGPT2 | N/A* | |||||
| GReaT Tabby MH DGPT2 | * | |||||
| GTT Base DGPT2 | ||||||
| GTT Tabby MH DGPT2 |
| Diabetes | Travel | Adult | Abalone | Rainfall | House | |
| CTGAN | ||||||
| TVAE | ||||||
| Tab-DDPM | ||||||
| Plain Base DGPT2 | ||||||
| Plain Tabby MH DGPT2 | ||||||
| GReaT Base DGPT2 | N/A* | |||||
| GReaT Tabby MH DGPT2 | * | |||||
| GTT Base DGPT2 | ||||||
| GTT Tabby MH DGPT2 |

### v2

@@ 3.2 Investigating the Choice of Base Model
Comparisons: We compare synthesis quality across LLMs of varying sizes. We consider 7 LLMs, listed in Table 3, evaluating Non-Tabby and MH versions of each. Each model is Plain-trained under conditions provided in Section 3.0, then sampled 500 times. Results are averaged across two runs. Llama models use LoRA [hu2021lora] on all linear transformer layers, with the LM head fully fine-tuned.
@@
We find that Tabby improves MLE or maintains upper-bound MLE for models, without necessarily increasing the cost of inference (Section 2.4). Although higher-parameter models are generally correlated with greater generative abilities, Figure 7 demonstrates that this is not always the case: Interestingly, we find that the Llama models (1.2B and 8B parameters each), have lower average MLE than smaller models. Tabby offers favorable performance improvements relative to the scaling curve and allows even small models to better outperform large, resource-intensive models.
@@ 3.3 Extending Tabby Beyond Tabular Data to General Structured Modalities
Comparisons: We plain-train non-Tabby and Tabby MH models on a JSON dataset of patients being evaluated for Glaucoma [manoj_glaucoma_2024]. Each datapoint has 10 features, organized in 3 groups: a group of 7 columns representing qualitative aspects of the optic nerve, a group of 2 columns corresponding to measurements between the optic nerve and eye, and a standalone feature for the diagnosis (examples in Box C). The binary classification target is inside the first group and assesses whether the optic nerve is thinning. As with tabular datasets in Section 3.1 and 3.2, we train downstream classifiers to predict the target variable and then present the resulting MLE.
We also consider the discrimination metric: Given equal numbers of real and synthetic samples, we measure the accuracy of a discrimination classifier that is trained to distinguish real versus synthetic datapoints. Because accuracy would indicate that the classifier is fully unable to distinguish real from synthetic, we report the accuracy’s distance from in Table 4 so that lower scores indicate higher-quality synthesis.
@@
| Non-Synthetic (Upper Bound) | ||
| CTGAN | ||
| Forest Diffusion | ||
| Tab-DDPM | ||
@@
Results: Table 4 demonstrates that Tabby MH improves MLE to parity with real data. Tabby MH’s lower discrimination score signifies this model’s samples are more realistic than non-Tabby samples.
@@ 3.4 Tracking the Adaptation to Individual Columns
We address our fourth claim by examining Tabby’s progress while fine-tuning on tabular data.
@@ 3.5 Discussion
We find that Tabby models synthesize high-quality data in a variety of settings. In particular, Plain-trained Tabby MH consistently outperforms all prior LLM-based approaches and is comparable to or better than diffusion-based approaches in most settings, despite Tabby enjoying greater flexibility under fewer assumptions than those made for diffusion models. The Tabby architecture modification allows LLMs to better model both univariate column distributions and multivariate relationships across columns.
Unusually, we find that the baseline Plain training technique with Distilled-GPT2 performs quite well on several standard evaluation datasets. The high performance of the Plain training technique compared to prior LLM works on tabular synthesis, which are more complex, is surprising. Notably, GPT-2 and related models can even outperform much larger models such as Llama on these benchmarks. A plausible explanation is that tabular generative ability may depend on the proportion of parameters allocated to the LM head relative to the rest of the model—about 33% in GPT-2 versus roughly 7% in Llama-8B—suggesting that heavier output heads may better capture the structured dependencies characteristic of tabular data.
As of this writing, the Adult, House and Diabetes datasets have become quite prevalent for tabular synthesis evaluation. We hope that future research will build off of our evaluation setup by continuing to include more diverse and challenging tabular datasets, along with extensions to other structured modalities.
Limitations: We highlight two additional priorities for future work:
Privacy preservation is important for trainsets that contain sensitive data, such as patient medical information. While Tabby’s privacy preservation is similar to prior works (Appendix E), a method with strong formal privacy guarantees is an important next step for privacy-critical applications.
Computationally-constrained environments or tasks with particularly large datasets may require deep learning approaches that are specifically designed with efficiency in mind. The Plain training method and insights from Section 3.2 may be useful towards this priority, but we leave further experimentation and the efficient implementations detailed in Section 2.4 to future work.
@@ 4 Related Work
Tabular data has played a central role in machine learning since the field’s early days. In particular, decision trees and relatives [song_decision_2015] are well-adapted to table classification or regression. Table synthesis is a growing area, though frequently overlooked in favor of image and text synthesis.
Classical synthesis: Classic machine learning models, such as random forest models or Bayesian networks, may be used to synthesize tables [reiter2005using, zhang2017privbayes], but are limited in the data types and distributions that may be represented.
Generative Adversarial Networks (GANs): Many tabular synthesis methods rely on GANs [goodfellow2014generative, xu_modeling_2019], but have encountered several limitations. In particular, distributions of ordinal columns are frequently imbalanced, leading GANs to undesirable phenomena such as mode collapse. Continuous columns may possess multiple modes and complex distributions, which GANs also struggle to capture [xu_modeling_2019].
Diffusion Models: Forest Diffusion [jolicoeur-martineau_generating_2024] and Tab-DDPM [kotelnikov_tabddpm_2022] are state-of-the-art table synthesis approaches based on the diffusion model. Both show top performance on many standard tabular metrics and are reliable for certain applications. Unfortunately, this performance is achieved by strong assumptions on the nature of the data space–for instance, numeric target variables may only assume integer values (see Figure 4) and diffusion models are unable to model non-categorical string columns such as addresses or telephone numbers. The ability to reach table synthesis performance comparable to that of diffusion models, but with fewer assumptions, is as an area of active research.
LLMs: A small, but growing, body of work has applied LLMs’ flexible modeling abilities to tables. GReaT [borisov_language_2022] is a method to convert tabular data into a sentence format compatible with LLMs, then “shuffling" the order in which columns occur for each row to improve the modeling of inter-column dependencies. TapTap [zhang_generative_2023] pretrains LMs on a variety of tabular data before fine-tuning on a downstream table synthesis task, while Tabula [zhao_tabula_2023] explores methods of preprocessing the training data to decrease sequence length. Other LLM-based works have adapted these advances to relational tables [solatorio_realtabformer_2023], or used the emergent abilities of very large models such as GPT-4 to generate synthetic data using In-Context Learning in place of fine-tuning [seedat_curated_2024]. Many of these works can be used in concert with Tabby, as demonstrated in Section 3, and they offer the additional advantage over other architectures of enabling pretrained tabular models to adapt effectively to new datasets.
MoE Architectures: The key innovation of Tabby is the application of Gated Mixture of Expert (MoE) layers [gatedmoe1, gatedmoe2] for LLM table synthesis. MoE layers have enjoyed utility in multitask [moemultitask1, moemultitask2] and multimodal learning [moemultimodal1, moemultimodal3], by creating sets of model parameters dedicated to a specific task.
@@ 5 Conclusion
Tabby is an MOE-based architecture modification that allows LLMs to generate realistic tabular data. Tabby reaches MLE parity with real data in datasets. We hope this promising performance spurs future work on architecture modifications that allow LLMs to represent structured data.
Broader Impact Statement
Tabby is an improvement in the realism of synthetic tabular data, with extensions to non-tabular structured data. As such, Tabby (and its downstream applications) will have positive impacts on tasks that require synthetic structured data, such as low-data downstream tasks or privacy-critical tasks. However, Tabby may also be useful for negative downstream applications, such as the falsification of data—a drawback common to many advancements in the realism of generative modeling.
@@ Appendix A Terminology Glossary
For convenience, in addition to definitions within the main text, we list and define the most frequently-used terms and abbreviations in our paper here:
DGPT2: Distilled-GPT2 [gpt2].
@@
GReaT [borisov_language_2022]: The landmark work on fine-tuning pre-existing LLMs to synthesize tabular data by encoding datapoints as text. Similar to Plain training, but includes train-time complications such as shuffling the order in which columns are encoded and sample-time complications such as the inability to generate label values that do not occur in the training dataset. Discussed in-detail in Section 4.
GReaT+Tabula (GT): The combination of GReaT training plus Tabula [zhao_tabula_2023] data encoding; see Section 4.
GReaT+TapTap+Tabula (GTT): The combination of GReaT training plus Tabula encoding and TapTap [zhang_generative_2023] pre-training on tabular data (which is performed after the LLM is pre-trained on text data).
Low Rank Adapters (LoRA): Parameter-efficient training method from [hu2021lora].
@@
Non-Synthetic (Upper Bound): Used for the MLE metric, this score represents the performance of a downstream classifier trained on real, instead of synthetic, data. See Appendix D for details.
@@ Appendix B Additional dataset information
| Magic | https://archive.ics.uci.edu/dataset/159/magic+gamma+telescope |
| Shoppers | https://archive.ics.uci.edu/dataset/468/online+shoppers+purchasing+intention+dataset |
@@
Diabetes [diabetes] contains medical information on female hospital patients, including age, number of pregnancies and skin thickness. Downstream models learn to predict whether a given patient suffers from diabetes. Apart from the label, all dataset columns are numerical, with some columns taking only integer values, while others are floats.
Travel [tejashvi_tour_2023] was collected by a travel agency wishing to predict customer churn. With the binary variable churn as the target, features include whether the traveler booked a hotel, frequent flyer status and traveler age. While most features are categorical, there are two numerical columns: traveler age and the number of times that the customer has used the travel agency in the past.
Adult [adult] is a dataset commonly used to benchmark tabular classification algorithms. Each row contains basic information on one American adult, such as their age, years of education and marital status. For each adult, the downstream task is to predict whether their annual income is above or below . The features are a mix of categorical and numerical columns, with each numerical column taking only integer values.
The Magic [magic_gamma_telescope_159] dataset consists of Monte Carlo–simulated observations from the Cherenkov gamma telescope, designed to differentiate high-energy gamma-ray events from background hadronic showers. Each row represents a single event characterized by 10 continuous features derived from Cherenkov image parameters, such as length, width, and asymmetry. The target variable is binary, indicating whether the event originated from a gamma signal or a hadronic background, making the dataset suitable for evaluating classification performance on physics-based data.
The Shoppers [online_shoppers_purchasing_intention_dataset_468] dataset contains session-level data from an e-commerce website over a one-year period. Each record describes a single visit using 18 behavioral and contextual features, including page visit counts and durations, bounce and exit rates, visitor type, month, and weekend indicator. The target variable is binary, indicating whether the session resulted in a purchase, making the dataset suitable for evaluating classification methods on imbalanced user-behavior data.
Our first regression dataset is Abalone [abalone], which records the basic measurements of abalones, such weight and height. The target variable is the abalone’s age.
The Rainfall [zaman_machine_2018] dataset, while challenging to many LLM-based synthesis methods, contains only four columns which record historical weather data in Bangladesh. Its target variable is the amount of rainfall recorded, and the features are the year, month and weather station location.
House [house] is a standard regression dataset. Each row represents a block of houses in California during the 1990 census. The dataset records the number of households residing in the block, the block’s median building age, average number of bedrooms, and other basic information. The dataset’s target column is the block’s median house value, which is numerical and allows us to assess Tabby’s synthetic data in a regression task.
Glaucoma [manoj_glaucoma_2024] dataset consists of JSON records describing ophthalmic patients under consideration for a glaucoma diagnosis. Each record contains various qualitative and quantitative information about the eye, as demonstrated by the examples in Box C.
@@ Appendix D Details on Experimental Setup
Discrimination [disc] quantifies the degree to which the generative model introduces spurious correlations or other patterns that differentiate synthetic from real data. Given the real training dataset and a synthetic dataset , we sample the same number of rows from each. Next, we train a random forest classifier to discriminate between real and synthetic examples. Highest-quality synthetic data will result in discrimination accuracy, indicating that is unable to distinguish between and . For this reason, our reported discrimination scores are calculated as the absolute difference between and the accuracy of discriminator . Accordingly, lower discrimination scores represent better performance.
Distance to Closest Record: (DCR)
Distance to Closest Record (DCR) [lautrup_syntheval_2024] quantifies the distance between each synthetic datapoint and its most-similar example in the training set . In addition to synthesis quality, this metric is an indication of the degree to which the model memorizes samples during training. Specifically, for each synthetic example , we compute its distance to every training point (using L0 distance for categorical columns and L1 distance for numerical columns) and take the smallest of these distances. The overall DCR is then reported as the average of these minimum distances across all synthetic examples in SS. Lower DCR is associated with high-quality synthesis, but a DCR score of 0 implies that most synthetic examples are merely copies of training dataset points memorized during training. As such, we consider the best DCR to be the lowest nonzero score.
@@ E.1 Additional Metrics for Main Results Tables
For the experiment presented in Section 3.1, we include five additional metrics:
1.
Discrimination (Table 6) works similarly to a GAN’s discriminator [goodfellow2014generative]: the downstream discriminator model receives equal numbers of real and of synthetic datapoints, then learns to distinguish between them. The more difficulty that the discriminator model encounters in distinguishing between real and synthetic, the more we can say that the real dataset’s patterns are preserved within the synthetic examples.
2.
Distance to Closest Record (DCR, Table 7) measures how far the average synthetic datapoint lies from its nearest non-synthetic datapoint. If DCR equals zero, it indicates that the model has memorized its trainset, while a very large DCR indicates that the model is not preserving the trainset’s patterns very well: small, non-zero, DCR scores are ideal.
3.
Shape [shi_tabdiff_2025] (Table 8) measures, for each individual column, how well the synthetic column matches the real column’s distribution. We follow the SDMetrics library’s implementation, which uses the Kolmogorov-Smirnov statistic for categorical and the complement of Total Variation Distance for numerical columns, then averages the distances across all columns to report a final summary number.
4.
Trend [shi_tabdiff_2025] (Table 9)—which is often used in conjunction with Shape—measures relationships between columns, by quantifying the degree to which these relationships in the real data are preserved by the synthetic data. We again use the SDMetrics library’s implementation, which measures Pearson correlation for relationships between numerical columns and complement of Total Variation Distance for relationships between categorical columns or one categorical and one numerical column (which is first binned to discretize the values).
5.
Wasserstein Distance (Table 10) computes the distance between the real and synthetic datasets’ numeric columns. This metric allows us to capture the similarity of relationships among all numeric columns, as opposed to the pairwise interactions measured by the Trend metric.
6.
Membership Inference Attacks [yao2025dcrdelusionmeasuringprivacy] (MIA, Table 11) assess the likelihood of a synthesis method to leak the contents of its training data. We use the implementation from [german2025tabmia], which evaluates the performance of 50 different attack strategies on each synthesis method. We focus on the Abalone dataset and aggregate results by providing the maximum AUROC, mean AUROC and maximum accuracy across the 50 attacks for Tabby and other SOTA methods. For each metric, a lower number indicates a lower rate of trainset leakage and a better degree of privacy.
7.
Runtime (Table 12) is provided for Tabby and other SOTA methods. We report the time in minutes to train for one epoch and to produce 100 samples with one NVIDIA RTX A6000. Each method is run using the out-of-the-box implementation available online.
@@
These metrics largely corroborate our findings in Section 3.1. In particular, Plain Tabby MH’s low DCR and discrimination scores indicate that this model’s synthetic data closely resembles that of real data. Additionally, the DCR scores are small but nonzero, which indicates that the model is generating novel datapoints rather than simply repeating datapoints memorized during training.
Additionally, we provide results for the HARMONIC method [wang_harmonic_2024] in Table 13.
| Diabetes | Travel | Adult | Magic | Shoppers | Abalone | Rainfall | House | |
| CTGAN | ||||||||
| CTGAN+ | ||||||||
| TVAE | ||||||||
| CLLM | N/A* | |||||||
| TabSyn | ||||||||
| TabDiff | ||||||||
| Forest Diffusion | ||||||||
| Tab-DDPM | ||||||||
| Plain Base | ||||||||
| Plain Tabby MH | ||||||||
| GReaT Base | N/A* | |||||||
| GReaT Tabby MH | * | |||||||
| GTT Base | ||||||||
| GTT Tabby MH |
| Diabetes | Travel | Adult | Magic | Shoppers | Abalone | Rainfall | House | |
| CTGAN | ||||||||
| CTGAN+ | ||||||||
| TVAE | ||||||||
| CLLM | N/A* | |||||||
| TabSyn | ||||||||
| TabDiff | ||||||||
| Forest Diffusion | ||||||||
| Tab-DDPM | ||||||||
| Plain Base | ||||||||
| Plain Tabby MH | ||||||||
| GReaT Base | N/A* | |||||||
| GReaT Tabby MH | * | |||||||
| GTT Base | ||||||||
| GTT Tabby MH |
| Diabetes | Travel | Adult | Magic | Shoppers | Abalone | Rainfall | House | |
| TabSyn | ||||||||
| TabDiff | ||||||||
| Forest Diffusion | ||||||||
| Tab-DDPM | ||||||||
| Plain Base | ||||||||
| Plain Tabby MH | ||||||||
| GReaT Base | ||||||||
| GReaT Tabby MH | ||||||||
| GTT Base | ||||||||
| GTT Tabby MH |
| Diabetes | Travel | Adult | Magic | Shopping | Abalone | Rainfall | House | |
| TabSyn | ||||||||
| TabDiff | ||||||||
| Forest Diffusion | ||||||||
| Tab-DDPM | ||||||||
| Plain Base | ||||||||
| Plain Tabby MH | ||||||||
| GReaT Base | N/A* | |||||||
| GReaT Tabby MH | ||||||||
| GTT Base | ||||||||
| GTT Tabby MH |
| Diabetes | Travel | Adult | Magic | Shopping | Abalone | Rainfall | House | |
| TabSyn | ||||||||
| TabDiff | ||||||||
| Forest Diffusion | ||||||||
| Tab-DDPM | ||||||||
| Plain Base | ||||||||
| Plain Tabby MH | ||||||||
| GReaT Base | N/A* | |||||||
| GReaT Tabby MH | ||||||||
| GTT Base | ||||||||
| GTT Tabby MH |
| TabDiff | CTGAN+ |
CLLM | HARMONIC | |||||||
| Max AUROC () | |||||||||||||
| Mean AUROC () | |||||||||||||
| Max Acc. () |
HARMONIC | CTGAN | CGAN+ | CLLM | TabDiff | TabSyn | |||||||
| Training | 2:05 | 2:01 | 2:03 | 24:25 | 0:01 | 0:01 | N/A | 0:02 | 0:02 | ||||||
| Synthesis | 0:10 | 0:10 | 0:11 | 7:30 | 0:01 | 0:01 | 0:04 | 0:00 | 0:00 |
@@
| MLE | ||||||
| Discrimination | ||||||
| DCR | ||||||
| Shape | ||||||
| Trend | ||||||
| Wasserstein |